OpenStudy (anonymous):

Solve the equation for x: x+4 = sqrt(8-x)

OpenStudy (jhannybean):

Square both sides of the equation first.

OpenStudy (alexandervonhumboldt2):

(x+4)^2=8-x

OpenStudy (alexandervonhumboldt2):

x^2+8x+16=8-x

OpenStudy (jhannybean):

\[(x+4)^2 = (\sqrt{8-x})^2\]

OpenStudy (alexandervonhumboldt2):

x^2+9x+8=0

OpenStudy (alexandervonhumboldt2):

do you know quadratic formula? @nlom

OpenStudy (jhannybean):

You don't need it.

OpenStudy (alexandervonhumboldt2):

hmm yeah

OpenStudy (jhannybean):

Fidn two numebrs that add to give you 9, and multiply to give you 8.

OpenStudy (anonymous):

\[x+4=\sqrt{8-x}\] \[8-x \ge0,8\ge x ~or~x \le 8.\] squaring \[\left( x+4 \right)^2=8-x\] \[x^2+8x+16-8+x=0\] \[x^2+9x+8=0\] find x by quadratic formula or by making factors.

OpenStudy (zale101):

@nlon , you need to participate on your post.

OpenStudy (zale101):

@nlom

OpenStudy (alexandervonhumboldt2):

(-9+-sqrt(81-32))/2

OpenStudy (alexandervonhumboldt2):

(-9+-7)/2

OpenStudy (alexandervonhumboldt2):

tell me what you get

OpenStudy (anonymous):

From squaring both sides I got x^2 + 16 = 8-x ?

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (alexandervonhumboldt2):

(a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

Thanks!

OpenStudy (alexandervonhumboldt2):

np